OpenStudy (anonymous):

Linda invests $5000 into an account that pays 11% compounded quarterly. How much interest will she get after 5 years?

OpenStudy (nowhereman):

Is her gain reinvested instantly or paid out?

OpenStudy (anonymous):

i dont no. thats the exact question. i copied n pasted it

OpenStudy (nowhereman):

http://en.wikipedia.org/wiki/Compound_interest It is revinvested. That means you have to calculate \[5000\cdot 1.11^{4\cdot 5}\]

OpenStudy (anonymous):

wait how do i write 11% ??? 0.11 or 1.11 ??? ... and i dont no how to do that equation on a calculator

OpenStudy (anonymous):

the final answer is 35311.56$

OpenStudy (anonymous):

you do it like this:

OpenStudy (nowhereman):

11% = 0.11 but the original sum remains there too, so you have 1 + 11% = 1.11

OpenStudy (anonymous):

\[5000(1+(11/100))^{5*4}\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

this is how i wrote it .. .A = 5000 (1 = 0.11/4) 4x5

OpenStudy (anonymous):

no, that is wrong

OpenStudy (anonymous):

i cant seem to figure that out on a calculator, u told me the answer but i cant get it on the calculator

OpenStudy (nowhereman):

Is it 11% anually or quarterly?

OpenStudy (anonymous):

i meant to write 1 + 0.11

OpenStudy (anonymous):

quarterely

OpenStudy (anonymous):

ok, the full formula for compound interest is this \[A=P(1+rate)^{number of years * timez per year}\] so the timezperyear is 4 in your case as there are 4 quoters in a year

OpenStudy (nowhereman):

Ok, so that /4 was a mistake of yours. Well, what kind of functions does your calculator have? Is there a button that says "a^b" or maybe "x^y" ? Or otherwise there maybe buttons for "exp" and "ln"

OpenStudy (anonymous):

umm im doing it on my cellphone calculator so no

OpenStudy (anonymous):

in a calculator it would look like A=P(1+rate)^(numberofyears∗timezperyear)

OpenStudy (anonymous):

damn, you should get a casio fx series

OpenStudy (nowhereman):

you can also do it on google or wolframalpha

OpenStudy (anonymous):

sooo how do i write down the equation

OpenStudy (anonymous):

it can solve integrals, derivatives, logarithms, anything

OpenStudy (anonymous):

like this: A = 5000(1+(11/100))^5∗4

OpenStudy (anonymous):

which is: = 5000(1.11)^20

OpenStudy (anonymous):

which is A = 5000 * 8.062311536 = 40311.55768 so now to get the compound interest subtract the 5000 from the answer, CI=40311.55768-5000 =35311.56

OpenStudy (anonymous):

ok thank uu

OpenStudy (anonymous):

one more?? Using the rule of 72, how long will it take to double your money if you invest in an account earning 3% interest? Give units of time (days, weeks, months, or years)

OpenStudy (anonymous):

24 years

OpenStudy (anonymous):

ok i got it right thx :)

OpenStudy (anonymous):

you just divide 72 by the percentage

OpenStudy (anonymous):

wat does it mean tho by give units of time (days,weeks, months, years???

OpenStudy (anonymous):

or, if the question asks you how long it takes for the money to halve, divide 72 by percentage :)

OpenStudy (anonymous):

it is there because the answer is not always in years, for example, if a question asks you how long it takes to double your money if you invest in an account earning 80% interest. It is IMPOSSIBLE in real life, but some questons might ask that, so, 72/80=0.9, which is 0.9 year, which is 365/10*9=328.5=329days

OpenStudy (anonymous):

get it? but most of the times the answer is in years, so dont worry

OpenStudy (anonymous):

can you see if these are right for me please. 6. Find the simple interest on $500 for 3 years at 7%. is it like this 500 x 0.07 x 3 = $105 7. Find the simple interest on $800 at 8.75% for 100 days. is it like this $800 x .0875 x 100/365 = $19.18

OpenStudy (anonymous):

6. is right 7. is right

OpenStudy (anonymous):

sweet thanks and is this one right ... (last one) How much money would you have to invest to make $300 within 2 years if the interest rate was 4.5%? is the answer $3333.33

OpenStudy (anonymous):

i think like 3 diff people gave me different answers, so im so confused!!

OpenStudy (anonymous):

thats right